OpenStudy (kendricklamar2014):

Write the value of the expression: 3^3/3^6

OpenStudy (kendricklamar2014):

@AccessDenied

OpenStudy (kendricklamar2014):

@agent0smith

OpenStudy (anonymous):

3^3=27 and 3^6=729. So 27/729. Now you simplify.

OpenStudy (solomonzelman):

subtract exponents.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 3^3/3^6=3^{3-6}=3^{-3}=1/3^3=1/27 }\)

OpenStudy (kendricklamar2014):

thanks

OpenStudy (kendricklamar2014):

How would u do this one? 3^4/3^4

OpenStudy (kendricklamar2014):

@SolomonZelman

OpenStudy (anonymous):

81/81 which is equal to 1.

OpenStudy (anonymous):

3*3*3*3=81

OpenStudy (solomonzelman):

a/a =1 now imagine that 3^4 is "a"

OpenStudy (kendricklamar2014):

ok

OpenStudy (kendricklamar2014):

thanks